OpenStudy (anonymous):

simplify 49sqrt14 over 21sqrt3

OpenStudy (anonymous):

no no simplify

OpenStudy (anonymous):

rationalize denominator and simplify

OpenStudy (anonymous):

\[\frac{49\sqrt{14}}{21\sqrt3}=\frac{49\sqrt{14}\cdot\sqrt{3}}{21\sqrt3\cdot{3}}=\frac{49\sqrt{42}}{62}\]

OpenStudy (anonymous):

sorry missed square in denominator should be \[21\sqrt3\cdot\sqrt3\]

OpenStudy (anonymous):

hold on let me put in equation

OpenStudy (anonymous):

\[49\sqrt{14} \over 21\sqrt{3}\]

jhonyy9 (jhonyy9):

7sqrt14 -------- 3sqrt3